Using mean value theorem find the √11
1
Expert's answer
2020-03-04T17:59:06-0500

Let "y=f(x)=\sqrt{x}."

The mean value theorem states that


"f(b) - f(a) = f'(c)(b-a), \quad a<c<b"

Let "b" is near "a" then we can write "\Delta x=b-a" and rewrite the formula as


"\Delta y= f'(c)\Delta x"

If "\Delta x \to 0," then we obtain


"dy=f'(x)dx"

Thus, "f(x+\Delta x) - f(x) \approx f'(x)dx."

Therefore, "f(x+\Delta x) \approx f(x) + f'(x)dx."

In this case,

"x+\Delta x=11, x = 9, \Delta x=dx=2, \text{ and } f'(x) = \frac{1}{2\sqrt{x}}"

"f(11)= \sqrt{11} \approx \sqrt{9} + \frac{1}{2\sqrt{9}}\cdot 2 = 3 + \frac{1}{3} \approx 3.3333"


Answer: "\sqrt{11} \approx 3.3333."